OpenStudy (anonymous):

Show all work in dividing y^2-y-12/y^2-6y+8 by y^2+2y-15/y^2-&y+10. Use complete sentences to explain how to simplify this expression.

OpenStudy (dan815):

y^2-y-12/y^2-6y+8 * y^2-7y+10/ y^2+2y-15 (y - 4) (y +3)/ (y-4)(y-2) * (y -5)(y-2) / (y +5)(y-3) = (y +3) (y -5)/(y+5)(y-3) = y^2-2y - 15 / y^2+2y-15

OpenStudy (dan815):

just factor factor factor